We wrap a light, flexible cable around a solid cylinder with mass 0.68 kg  and radius 0.17 m  . The cylinder rotates with negligible friction about a stationary horizontal axis. We tie the free end of the cable to a block of mass 1.42 kg  and release the object with no initial velocity at a distance 1.63 m  above the floor. As the block falls, the cable unwinds without stretching or slipping, turning the cylinder. Suppose the falling mass is made of ideal rubber, so that no mechanical energy is lost when the mass hits the ground.

If the cylinder is originally not rotating and the block is released from rest at a height 1.63 m  above the ground, to what height will this mass rebound if it bounces straight back up from the floor?   
Answer #1

given
M = 0.68 kg
R = 0.17 m
m = 1.42 kg
h = 1.63 m

let v is the speed of hanging block and w is the angular speed of the cyllinder when the block hits the ground.

Apply conservation of energy

gain in kinetic energy of hanging block and cyllinder = loss of potential energy of hanging block

(1/2)*m*v^2 + (1/2)*I*w^2 = m*g*h

(1/2)*m*v^2 + (1/2)*(1/2)*M*R^2*w^2 = m*g*h

(1/2)*m*v^2 + (1/4)*M*v^2 = m*g*h (since v = R*w)

v^2*(m/2 + M/4) = m*g*h

v = sqrt(m*g*h/(m/2 + M/4))

= sqrt(1.42*9.8*1.63/(1.42/2 + 0.68/4) )

= 5.08 m/s

height raised by the block aabove the ground, H = v^2/(2*g)

= 5.08^2/(2*9.8)

= 1.32 m <<<<<<<<<<<<<<<<<---------------------------Answer